OpenStudy (anonymous):

Your plan is to rent a car and drive through the major European capitals. By consulting a map you estimate that you will cover a total distance of 5000 km. You select a rental package that includes a car with an average consumption of 6.00liters of fuel per 100 km. Considering that in Europe the average fuel cost is 1.063 euros per liter, how much (in US dollars) will you spend in fuel on your trip?

OpenStudy (agent47):

6-100 x-5000 x=5000*6/100=50*6=300 <= amount of fuel that you will spend. Multiply this amount by 1.063 euros

OpenStudy (agent47):

BAM, then a miracle occurs and u get the answer